OpenStudy (anonymous):

Find the unknown side length for the triangle. Area = 18 ft² A. 1.5 ft B. 3ft C. 6ft D. 9 ft Medal and fan Please help !!!

OpenStudy (anonymous):



OpenStudy (anonymous):

CAN SOMEONE HELP ME :)

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@J-bird

OpenStudy (anonymous):

sorry im here now

OpenStudy (anonymous):

Please tell me i am helping my cousin with her homework and i don't remember this when i was in middle school

OpenStudy (anonymous):

She is sick

OpenStudy (anonymous):

im trying the best i can myself my answer is A again, 1.5

OpenStudy (anonymous):

@JFraser @TTP @LunyMoony

OpenStudy (anonymous):

@mathmath333 @helpout3

OpenStudy (anonymous):

MEDAL AND FAN WHOEVER ANSWERS AND I GET THIS CORRECT THEY WILL GET MEDAL AND FAN PLEASE HURRAY

OpenStudy (lunymoony):

The at the bottom side for an answer?

OpenStudy (anonymous):

@anonymous_user

OpenStudy (anonymous):



OpenStudy (anonymous):

The area is 18ft please help

OpenStudy (anonymous):

@wizfan

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (anonymous):

SOMEONE HELP PLEASE MEDAL AND FAN

OpenStudy (anonymous):

@brainbuzzed5646 @theEric @hhmh4 @Hauntedwoodsgal @theEric

OpenStudy (lunymoony):

C, I believe

OpenStudy (anonymous):

C thx you get a medal and fan

OpenStudy (lunymoony):

np

OpenStudy (anonymous):

Thx sooo much i am helping my cousin with her homework and i don't remember stuff like this from middle schoool

OpenStudy (lunymoony):

Ok, Cool

OpenStudy (mathmath333):

i think the information seems insufficient here

OpenStudy (anonymous):

u are too late please help me now k

OpenStudy (mathmath333):

if it is right angled triangle here then the unknown side would be 3ft ie B

OpenStudy (anonymous):

Which equation can be used to solve the problem? How many 8-gallon cans are needed to hold a total of 120 gallons of water? A. c/8=120 B. 8/c =120 C. 8+c=120 D. 8c=120

OpenStudy (anonymous):

@mathmath333 @bibby

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (lunymoony):

This one seems too simple to be believed, it is D. 120 gallons of water divided by 8 gallon cans = the answer, or 15